OpenStudy (anonymous):

15+(21/7-10 is it 8?

OpenStudy (anonymous):

Is the problem 15+(21/7-1)?

OpenStudy (anonymous):

Ok sorry i got it now let me see

OpenStudy (anonymous):

15+ 21 7 −10 =15+3−10 =15+−7 =8

OpenStudy (anonymous):

So yes you are right

OpenStudy (anonymous):

:)